OpenStudy (ell):

Solve each equation by factoring: n^2 + 3n - 12 = 6

OpenStudy (kenljw):

n^2 + 3n - 12 = 6 n^2 + 3n-18 =0 at this point you can use the quadratic eq or recognizie (n+6)(n-3)

OpenStudy (kenljw):

The way you recognize (n+a)(n+b) = n^2 + (a+b)n +ab

OpenStudy (ell):

i get the first part, but now im lost.. sorry!

OpenStudy (kenljw):

(n+a)(n+b) = n^2 + (a+b)n +ab use distributive law and multiple out

OpenStudy (anonymous):

n^2 + 3n-18 =0 basically focus on left side what numbers multiply to get 18 but sum up to 3. and that is 6 and -3

OpenStudy (ell):

okay! thanks @timo86m that helped alot:P so the answers are (6,-3) ?

OpenStudy (ell):

thanx! :)

OpenStudy (anonymous):

(n+6)(n-3)= n^2 + 3n-18 = 0 since (n+6)(n-3)=0 is easier to solve than n^2 + 3n-18 = 0 Just solve (n+6)(n-3)=0 n=-6 and n=3

OpenStudy (ell):

Thank you so much! too the both of you! you guys are such a big help! much appreciated! <3 :)

OpenStudy (kenljw):

(n+a)(n+b) = n^2 + (a+b)n +ab use distributive law and multiple out Therefore and Quadric eq of the form n^2 + Bn + C B = a+b C=ab

OpenStudy (ell):

thanx! GBU!